# RGNMR: A Gauss-Newton method for robust matrix completion with theoretical guarantees

### v1

@@ Appendix B Additional simulations
Figure 5: Failure probability (±$\pm 1.96$SE) of RMC methods as a function of the fraction of corrupted entries $\alpha$. (left) Oversampling ratio of $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=8$; (right) Oversampling ratio of $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=12$. The underlying matrix $L^{*}$ if of size $3200\times 400$ has a rank $r=5$ and a condition number $\kappa=2$. Each point corresponds to 100 independent realizations.
Figure 6: Results of RMC methods as a function of the oversampling ratio under non uniform sampling. (left) Median rel-RMSE ; (right) Failure probability defined as $\mathbb{P}(\texttt{rel-RMSE}>10^{-3})$. The underlying matrix $L^{*}$ is of size $1000\times 1000$ has rank $r=5$ and condition number $\kappa=2$. The fraction of corrupted entries is $\alpha=5\%$. Each point corresponds to 50 independent realizations.
@@
Figure 10: Median run time of RGNMR and RGNMR -BS for a matrix of size $n\times n$ as a function of $n$. The matrix has rank $r=5$ and a condition number $\kappa=2$. The oversampling ratio is $\rho=\frac{|\Omega|}{r(2n-r)}=6$ and the fraction of corruption is $\alpha=5\%$. Each point corresponds to 20 independent realizations. Both $x$ and $y$ axes are on a log scale and the dashed lines have a slope of 2.
@@
In many applications, the entries of the matrix are not sampled uniformly (Meka et al.,, 2009). Hence we made simulations where the observed entries followed a power law sampling scheme similar to (Meka et al.,, 2009), described in the following lines. Given an oversampling ratio $\rho$ we define $w=\rho\cdot r\cdot(n_{1}+n_{2}-r)$. We construct two sequences $(\tilde{p}_{1}\dots\tilde{p}_{n_{1}})$ and $(\tilde{q}_{1}\dots\tilde{q}_{n_{2}})$ such that $\tilde{p}_{i}=i^{-\frac{2}{3}}$, $\tilde{q}_{i}=i^{-\frac{2}{3}}.$ We then normalize them to construct two new sequences
@@
In Figure 7 we illustrate RGNMR stability to additive noise. In this simulation, all observed entries are corrupted by additive white Gaussian noise with known standard deviation $\sigma$, in addition to the few outliers entries. In this scenario inliers entries are corrupted with random noise. Note that our method for upper bounding $k^{*}$ is based on the assumption that the error in those entries is the result of rounding errors, see 2.1. To overcome this problem, when searching for an upper bound on $k^{*}$ we terminated RGNMR once $\frac{\|L_{t}-X\|_{F(\Omega\setminus\Lambda_{t})}}{\|X\|_{F(\Omega\setminus\Lambda_{t})}}\leq\sqrt{\sigma}$.
@@
Finally, In Figure 10 we illustrate that RGNMR scales well with the matrix size. We show that for a matrix of size $n\times n$ the runtime of RGNMR and RGNMR-BS increases quadratically with $n$.
@@
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ with $r$-th singular value $\sigma_{r}^{*}$. For large enough constants $c_{l},c_{e}$ the following holds: If $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}})\cap\mathcal{B}_{\mu}$ and for every $11\leq s\leq t+$ 1 , $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})$ , then $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})$.
@@
Let $X=L^{*}+S^{*}$ where $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ and the corruption matrix $S^{*}$ satisfies Assumption 3 for some known $10<\alpha<$ 1. Let the set of observed entries $\Omega$ follow Assumption 2 with $p\geq C\frac{\log n_{1}}{n_{2}}$ for some constant $C$. Let $L=UV^{\top}$ be an estimate of $L^{*}$ such that $(U,V)\in\mathcal{B}_{\mu}$. Suppose that the set of corrupted entries is estimated by $\Lambda=\text{support}\left(\mathcal{T}_{\gamma\alpha}\left(L-X,\Omega\right)\right)$ with an over removal factor of $1<\gamma\leq\frac{1}{\alpha}$. If there exists a factorization $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ of $L^{*}$ such that
@@
Eq. eq. 14 then follows immediately by the definition of $\mathcal{B}_{\textnormal{err}}$, see eq. 11.
Eq.(27) for $t=0$ follows from the assumption that the initialization is sufficiently accurate, balanced and has bounded row norms $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}})\cap\mathcal{B}_{\mu}$. Next we assume $(U_{t},V_{t})$ satisfies eq. 27 and prove that the updated matrices also satisfy this equation. Since at all intermediate steps $11\leq s\leq t+$ 1, $(U_{s},V_{s})$ are updated by eq. 10 , $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})\cap\mathcal{B}_{\mu}$. Hence by Lemma C.4, $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$.
@@
For the first part, to bound $\|L_{t+1}-L^{*}\|_{F({\Omega_{t})}}$ we apply Lemma C.5. To this end we first prove its conditions hold. In particular that the set $\mathcal{A}_{t}=\mathcal{B}^{*}\cap\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$ is non empty. Indeed this follows by applying Lemma C.1. Specifically, since $(U_{t},V_{t})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{2^{t}c_{e}\sqrt{\kappa}}\right)\cap\mathcal{B}_{\textnormal{bln}}\left(\frac{1}{c_{l}}\right)\cap\mathcal{B}_{\mu}$, for a large enough $c_{e}$, $\left(U_{t},V_{t}\right)$ satisfy the conditions of Lemma C.1 with $\epsilon=\frac{1}{20}$. Consequently, w.p. at least $1-\frac{2}{n_{1}^{5}}$, there exists a pair $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that satisfy eqs. 19a and 19b with $U=U_{t},V=V_{t}$. Combining eq. 19a with the assumptions $\frac{25\sigma_{r}^{*}}{c_{e}^{2}\kappa}\leq\delta$ and $\|L_{t}-L^{*}\|\leq\frac{\sigma_{r}^{*}}{2^{t}c_{e}\sqrt{\kappa}}$ yields
@@
We upper bound the term $T_{1},T_{2}$ and $T_{3}$. The first term $T_{1}$ is bounded by eq. 19b of Lemma C.1 with $\epsilon=\frac{1}{20}$.
@@
and a similar inequality holds for $\Delta V_{t+1}$. Hence LABEL:{eq:KMO10_lemma71_consequence_assumption} holds with $t=\sigma_{r}^{*}$. Consequently by eq. 22 of C.3 with $\epsilon=\frac{1}{8\sqrt{c_{e}\kappa}}$, w.p. at least $1-\frac{2}{n_{1}^{5}}$
@@
In order to bound $T_{3}$ we invoke Lemma C.6. The condition of this Lemma, eq. 24, holds since we proved that $(U_{t},V_{t})$ satisfy eq. 19a of Lemma C.1. Since $\Lambda_{t}=\textnormal{support}\left(\mathcal{T}_{\theta}\left(L_{t}-X,\Omega\right)\right)$ we can employ Lemma C.6 which gives
@@
Since $(U_{t},V_{t})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{2^{t}c_{e}\sqrt{\kappa}})$, for a large enough $c_{e}$, $(U_{t},V_{t})$ satisfy the conditions of Lemma C.2 with $\epsilon=\frac{1}{20}$. Therefore by eq. 20 w.p. at least $1-\frac{3}{n_{1}^{3}}$,
@@
Inserting the bounds on $T_{1},T_{2}$ and $T_{3}$ from equations (29), (30) and (33) into eq. 28 gives
@@
To upper bound $\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Lambda_{t})}$ we apply Lemma C.7, with $(U_{t+1},V_{t+1})$ which indeed by their definition in eq. 10, belong to $B_{\mu}$, and with a specific factorization $(U^{*},V^{*})$ specified below. Specifically, we show that there exists a pair $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that is in the vicinity of $\left(U_{t+1},V_{t+1}\right)$. By Lemma C.4 $\left(U_{t+1},V_{t+1}\right)\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$. Therefore, for a large enough $c_{e}$ Lemma C.1 guarantees the existence of a factorization $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that also satisfy eq. 19a with $L=L_{t+1}=U_{t+1}V_{t+1}^{\top}$. Hence, by eq. 26 of Lemma C.7
@@
Combining eq. 19a and the assumptions $\alpha\leq\frac{1}{c_{\alpha}r\mu\kappa},\gamma\leq\sqrt{c_{\alpha}}$ gives
@@
Inserting eq. 37 and eq. 38 into eq. 35 yields
@@
By eq. 34 the right hand size is upper bounded by $\frac{\sigma_{r}^{*}}{2^{t+1}c_{e}\sqrt{\kappa}}$, which implies that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{2^{t+1}c_{e}\sqrt{\kappa}}\right)$. Since $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$ was proved at the beginning, $(U_{t+1},V_{t+1})$ satisfy eq. 27, which concludes the proof. ∎
@@
Let $X=L^{*}+S^{*}$ where $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu)$ with singular values $\sigma_{1}^{*}\geq\dots\geq\sigma_{r}^{*}$, and $S^{*}$ satisfies Assumption 3 for some known $10<\alpha<$ 1. Let $L=UV^{\top}$ be an estimate of $L^{*}$ constructed according to the following equation,
@@
Recall the definition of $S_{init}$, eq. 16. The following lemma bounds the operator norm of $Z=\text{b-SVD}_{r}{\big{[}}\frac{1}{p}\left(\mathcal{P}_{\Omega}(X)-S_{init}\right){\big{]}}$. Its proof appears in Appendix D.4.
@@
In the last step of the algorithm we apply the clipping operator $R_{\eta}$, see eq. 18. The following lemma , whose proof appears in Appendix D.5, state that under some conditions applying the clipping operator with the right $\eta$ reduces the Procrustes distance.
@@
[…]
@@
Denote $\xi=\left(\frac{16}{c_{\alpha}}+\frac{4c}{\sqrt{C}}\right)$. Inserting eq. 50 into eq. 48 gives
@@
Inserting the above bound into eq. 47 gives
@@
Next we prove that $(U_{0},V_{0})\in\mathcal{B}_{\mu}$, which requires to bound $\|U_{0}\|_{2,\infty}$ and $\|V_{0}\|_{2,\infty}$. Since $\begin{pmatrix}U_{0}\\ V_{0}\end{pmatrix}=\begin{pmatrix}R_{\eta_{1}}(U)\\ R_{\eta_{2}}(V)\end{pmatrix}$ by the definition, eq. 18, of the clipping operator $R_{\eta}$
@@
Since we proved that $\|UV^{\top}-L^{*}\|\leq\frac{\sigma_{1}^{*}}{4}$ we can apply Lemma C.11. Specifically inserting eq. 45 into the above inequality completes the proof as it gives
@@
Let $Z_{t}$ be the estimated factorization after $t$ steps as defined in Algorithm 2, and let $Z_{t+1}$ be the factorization at iteration $t+1$. Denote
@@
For large enough constants $c_{e},c_{l}$ the following holds. If $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}({1}/{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}({1}/{2c_{l}})\cap\mathcal{B}_{\mu}$ and $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})$ for every $11\leq s\leq t+$ 1 then
@@
We start by proving that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})$. By eq. 40 of Lemma C.8,
@@
By eq. 51 of Lemma D.1, for large enough $c_{e}$, for each iteration step $0\leq k\leq t$
@@
Inserting eq. 53 into eq. 52 gives,
@@
By the assumption $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{c_{e}\sqrt{\kappa}}\right)$ it follows that $\|U_{0}V_{0}^{\top}-L^{*}\|\leq\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}$. Inserting this inequality and eq. 55 into eq. 54 gives
@@
To upper bound the first term on the RHS of eq. 56 we use the following lemma.
@@
To upper bound the RHS of eq. 57 we employ the following two lemmas. The first lemma upper bounds the first term and the second lemma lower bounds the second term.
@@
To upper bound $\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})$ we insert eq. 59 and eq. 58 into eq. 57 which yields,
@@
Inserting eq. 61 into eq. 60 and using the fact that $\|S\|_{F({\Omega_{t}})}=\|S^{*}\|_{F(I_{t}^{\text{out}})}$ gives,
@@
Inserting eq. 63 and eq. 64 into eq. 62 gives,

### v2

@@ Appendix B Additional simulations
Figure 5: RGNMR performance compared to that of GNMR (Zilber and Nadler,, 2022) under a small fraction of corrupted entries. (left) Median rel-RMSE ; (right) Failure probability defined as $\mathbb{P}(\texttt{rel-RMSE}>10^{-3})$. The underlying matrix $L^{*}$ is of size $500\times 500$ has rank $r=5$ and condition number $\kappa=2$. The oversampling ratio is $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=12$ and the fraction of corrupted entries $\alpha$ varies. Each point corresponds to 50 independent realizations.
Figure 6: Failure probability ($\pm 1.96$SE) of RMC methods as a function of the fraction of corrupted entries $\alpha$. (left) Oversampling ratio of $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=8$; (right) Oversampling ratio of $\frac{|\Omega|}{r(n_{1}+n_{2}-r)}=12$. The underlying matrix $L^{*}$ if of size $3200\times 400$ has a rank $r=5$ and a condition number $\kappa=2$. Each point corresponds to 100 independent realizations.
Figure 7: Results of RMC methods as a function of the oversampling ratio under power law sampling. (left) Median rel-RMSE ; (right) Failure probability defined as $\mathbb{P}(\texttt{rel-RMSE}>10^{-3})$. The underlying matrix $L^{*}$ is of size $1000\times 1000$ has rank $r=5$ and condition number $\kappa=2$. The fraction of corrupted entries is $\alpha=5\%$. Each point corresponds to 50 independent realizations.
Figure 8: Results of RMC methods as a function of the oversampling ratio under diagonal-band pattern sampling. (left) Median rel-RMSE ; (right) Failure probability defined as $\mathbb{P}(\texttt{rel-RMSE}>10^{-3})$. The underlying matrix $L^{*}$ is of size $500\times 500$ has rank $r=5$ and condition number $\kappa=2$. The fraction of corrupted entries is $\alpha=5\%$. Each point corresponds to 50 independent realizations.
@@
Figure 10: Performance of RMC methods as a function of the condition number under both outliers and additive noise. (left) Median rel-RMSE; (right) Failure probability ($\pm 1.96$ SE). The underlying matrix $L^{*}$ is of size $3200\times 400$ has rank $r=5$, the fraction of corrupted entries is $\alpha=5\%$, the standard deviation of the white Gaussian noise is $\sigma=10^{-6}$ and the oversampling ratio is $\frac{|\Omega|}{r\cdot(n_{1}+n_{2}-r)}=12$. Each point corresponds to 100 independent realizations.
Figure 11: Performance of RMC methods as a function of the oversampling ratio under both outliers and additive noise. (left) Median rel-RMSE; (right) Failure probability ($\pm 1.96$ SE). The underlying matrix $L^{*}$ is of size $3200\times 400$ has rank $r=5$, the fraction of corrupted entries is $\alpha=5\%$ the standard deviation of the white Gaussian noise is $\sigma=10^{-6}$ and the condition number is $\kappa=2$. Each point corresponds to 100 independent realizations.
@@
Figure 14: Effect of the matrix size on RGNMR performance. Both $x$ and $y$ axes are on a log scale. (left) Median run time of RGNMR and RGNMR -BS for a matrix of size $n\times n$ as a function of $n$. The dashed lines have a slope of 2. (right) Median Error as a function of the matrix size. The matrix has rank $r=5$ and a condition number $\kappa=2$. The oversampling ratio is $\rho=\frac{|\Omega|}{r(2n-r)}=6$ and the fraction of corruption is $\alpha=5\%$. Each point corresponds to 20 independent realizations.
@@
In many applications, the entries of the matrix are not sampled uniformly (Meka et al.,, 2009; Okatani et al.,, 2011). Hence we made simulations where the observed entries followed a power law sampling scheme similar to Meka et al., (2009) and a diagonal-band pattern as in Okatani et al., (2011).
For the power law scheme, given an oversampling ratio $\rho$ we define $w=\rho\cdot r\cdot(n_{1}+n_{2}-r)$. We construct two sequences $(\tilde{p}_{1}\dots\tilde{p}_{n_{1}})$ and $(\tilde{q}_{1}\dots\tilde{q}_{n_{2}})$ such that $\tilde{p}_{i}=i^{-\frac{2}{3}}$, $\tilde{q}_{i}=i^{-\frac{2}{3}}.$ We then normalize them to construct two new sequences
@@
For the diagonal-band pattern, we generated $n\times n$ matrices of rank $r$ with a diagonal bandwidth of length $pr$ across different values of $p$. This results in approximately $(n+n)\cdot pr$ observed entries and therefore $p$ is approximately the oversampling ratio. As illustrated in Figure 8, though RGNMR requires a larger oversampling ratio than in the uniform pattern to succeed in this task, it stills outperforms other RMC methods.
@@
In Figures 9, 10 and 11 we illustrate that RGNMR still outperform most RMC methods even under additive noise. In these simulation, all observed entries are corrupted by additive white Gaussian noise with known standard deviation $\sigma$, in addition to the few outliers entries. In this scenario inliers entries are corrupted with random noise. Note that our method for upper bounding $k^{*}$ is based on the assumption that the error in those entries is the result of rounding errors, see 2.1. To overcome this problem, when searching for an upper bound on $k^{*}$ we terminated RGNMR once $\frac{\|L_{t}-X\|_{F(\Omega\setminus\Lambda_{t})}}{\|X\|_{F(\Omega\setminus\Lambda_{t})}}\leq\sqrt{\sigma}$.
@@
Finally, In Figure 14 we illustrate that RGNMR scales well with the matrix size. We show that for a matrix of size $n\times n$ the runtime of RGNMR and RGNMR-BS increases quadratically with $n$. In addition, we illustrate that even on large matrices RGNMR still requires a relatively small number of observed entries to succeed.
@@
Let $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ with $r$-th singular value $\sigma_{r}^{*}$. For large enough constants $c_{l},c_{e}$ the following holds: If $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}})\cap\mathcal{B}_{\mu}$ and for every $1\leq s\leq t+1$ , $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})$ , then $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})$.
@@
Let $X=L^{*}+S^{*}$ where $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu,\kappa)$ and the corruption matrix $S^{*}$ satisfies Assumption 3 for some known $0<\alpha<1$. Let the set of observed entries $\Omega$ follow Assumption 2 with $p\geq C\frac{\log n_{1}}{n_{2}}$ for some constant $C$. Let $L=UV^{\top}$ be an estimate of $L^{*}$ such that $(U,V)\in\mathcal{B}_{\mu}$. Suppose that the set of corrupted entries is estimated by $\Lambda=\text{support}\left(\mathcal{T}_{\gamma\alpha}\left(L-X,\Omega\right)\right)$ with an over removal factor of $1<\gamma\leq\frac{1}{\alpha}$. If there exists a factorization $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ of $L^{*}$ such that
@@
Eq. eq.˜14 then follows immediately by the definition of $\mathcal{B}_{\textnormal{err}}$, see eq.˜11.
Eq.(27) for $t=0$ follows from the assumption that the initialization is sufficiently accurate, balanced and has bounded row norms $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{2c_{l}})\cap\mathcal{B}_{\mu}$. Next we assume $(U_{t},V_{t})$ satisfies eq.˜27 and prove that the updated matrices also satisfy this equation. Since at all intermediate stpdf $1\leq s\leq t+1$, $(U_{s},V_{s})$ are updated by eq.˜10 , $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})\cap\mathcal{B}_{\mu}$. Hence by Lemma C.4, $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$.
@@
For the first part, to bound $\|L_{t+1}-L^{*}\|_{F({\Omega_{t})}}$ we apply Lemma C.5. To this end we first prove its conditions hold. In particular that the set $\mathcal{A}_{t}=\mathcal{B}^{*}\cap\mathcal{B}_{\mu}\cap\mathcal{C}\left(U_{t},V_{t},\frac{\delta}{4^{t+1}}\right)$ is non empty. Indeed this follows by applying Lemma C.1. Specifically, since $(U_{t},V_{t})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{2^{t}c_{e}\sqrt{\kappa}}\right)\cap\mathcal{B}_{\textnormal{bln}}\left(\frac{1}{c_{l}}\right)\cap\mathcal{B}_{\mu}$, for a large enough $c_{e}$, $\left(U_{t},V_{t}\right)$ satisfy the conditions of Lemma C.1 with $\epsilon=\frac{1}{20}$. Consequently, w.p. at least $1-\frac{2}{n_{1}^{5}}$, there exists a pair $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that satisfy eqs.˜19a and 19b with $U=U_{t},V=V_{t}$. Combining eq.˜19a with the assumptions $\frac{25\sigma_{r}^{*}}{c_{e}^{2}\kappa}\leq\delta$ and $\|L_{t}-L^{*}\|\leq\frac{\sigma_{r}^{*}}{2^{t}c_{e}\sqrt{\kappa}}$ yields
@@
We upper bound the term $T_{1},T_{2}$ and $T_{3}$. The first term $T_{1}$ is bounded by eq.˜19b of Lemma C.1 with $\epsilon=\frac{1}{20}$.
@@
and a similar inequality holds for $\Delta V_{t+1}$. Hence LABEL:{eq:KMO10_lemma71_consequence_assumption} holds with $t=\sigma_{r}^{*}$. Consequently by eq.˜22 of C.3 with $\epsilon=\frac{1}{8\sqrt{c_{e}\kappa}}$, w.p. at least $1-\frac{2}{n_{1}^{5}}$
@@
In order to bound $T_{3}$ we invoke Lemma C.6. The condition of this Lemma, eq.˜24, holds since we proved that $(U_{t},V_{t})$ satisfy eq.˜19a of Lemma C.1. Since $\Lambda_{t}=\textnormal{support}\left(\mathcal{T}_{\theta}\left(L_{t}-X,\Omega\right)\right)$ we can employ Lemma C.6 which gives
@@
Since $(U_{t},V_{t})\in\mathcal{B}_{\textnormal{err}}(\frac{1}{2^{t}c_{e}\sqrt{\kappa}})$, for a large enough $c_{e}$, $(U_{t},V_{t})$ satisfy the conditions of Lemma C.2 with $\epsilon=\frac{1}{20}$. Therefore by eq.˜20 w.p. at least $1-\frac{3}{n_{1}^{3}}$,
@@
Inserting the bounds on $T_{1},T_{2}$ and $T_{3}$ from equations (29), (30) and (33) into eq.˜28 gives
@@
To upper bound $\lVert L_{t+1}-L^{*}\rVert^{2}_{F(\Lambda_{t})}$ we apply Lemma C.7, with $(U_{t+1},V_{t+1})$ which indeed by their definition in eq.˜10, belong to $B_{\mu}$, and with a specific factorization $(U^{*},V^{*})$ specified below. Specifically, we show that there exists a pair $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that is in the vicinity of $\left(U_{t+1},V_{t+1}\right)$. By Lemma C.4 $\left(U_{t+1},V_{t+1}\right)\in\mathcal{B}_{\textnormal{err}}(\frac{13}{\sqrt{c_{e}}})\cap\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$. Therefore, for a large enough $c_{e}$ Lemma C.1 guarantees the existence of a factorization $(U^{*},V^{*})\in\mathcal{B}^{*}\cap\mathcal{B}_{\mu}$ that also satisfy eq.˜19a with $L=L_{t+1}=U_{t+1}V_{t+1}^{\top}$. Hence, by eq.˜26 of Lemma C.7
@@
Combining eq.˜19a and the assumptions $\alpha\leq\frac{1}{c_{\alpha}r\mu\kappa},\gamma\leq\sqrt{c_{\alpha}}$ gives
@@
Inserting eq.˜37 and eq.˜38 into eq.˜35 yields
@@
By eq.˜34 the right hand size is upper bounded by $\frac{\sigma_{r}^{*}}{2^{t+1}c_{e}\sqrt{\kappa}}$, which implies that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{2^{t+1}c_{e}\sqrt{\kappa}}\right)$. Since $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})\cap\mathcal{B}_{\mu}$ was proved at the beginning, $(U_{t+1},V_{t+1})$ satisfy eq.˜27, which concludes the proof. ∎
@@
Let $X=L^{*}+S^{*}$ where $L^{*}\in\mathcal{M}(n_{1},n_{2},r,\mu)$ with singular values $\sigma_{1}^{*}\geq\dots\geq\sigma_{r}^{*}$, and $S^{*}$ satisfies Assumption 3 for some known $0<\alpha<1$. Let $L=UV^{\top}$ be an estimate of $L^{*}$ constructed according to the following equation,
@@
Recall the definition of $S_{init}$, eq.˜16. The following lemma bounds the operator norm of $Z=\text{b-SVD}_{r}{\big[}\frac{1}{p}\left(\mathcal{P}_{\Omega}(X)-S_{init}\right){\big]}$. Its proof appears in Appendix D.4.
@@
In the last step of the algorithm we apply the clipping operator $R_{\eta}$, see eq.˜18. The following lemma , whose proof appears in Appendix D.5, state that under some conditions applying the clipping operator with the right $\eta$ reduces the Procrustes distance.
@@
Next we explicitly bound $d_{P}(Z,Z^{*})$. Since $Z$ is obtained using $\textnormal{b-SVD}_{r}$, $\|U^{\top}U-V^{\top}V\|_{F}=0$. Therefore by Lemma C.10
@@
Denote $\xi=\left(\frac{16}{c_{\alpha}}+\frac{4c}{\sqrt{C}}\right)$. Inserting eq.˜50 into eq.˜48 gives
@@
Inserting the above bound into eq.˜47 gives
@@
Next we prove that $(U_{0},V_{0})\in\mathcal{B}_{\mu}$, which requires to bound $\|U_{0}\|_{2,\infty}$ and $\|V_{0}\|_{2,\infty}$. Since $\begin{pmatrix}U_{0}\\ V_{0}\end{pmatrix}=\begin{pmatrix}R_{\eta_{1}}(U)\\ R_{\eta_{2}}(V)\end{pmatrix}$ by the definition, eq.˜18, of the clipping operator $R_{\eta}$
@@
Since we proved that $\|UV^{\top}-L^{*}\|\leq\frac{\sigma_{1}^{*}}{4}$ we can apply Lemma C.11. Specifically inserting eq.˜45 into the above inequality completes the proof as it gives
@@
Let $Z_{t}$ be the estimated factorization after $t$ stpdf as defined in Algorithm 2, and let $Z_{t+1}$ be the factorization at iteration $t+1$. Denote
@@
For large enough constants $c_{e},c_{l}$ the following holds. If $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}({1}/{c_{e}\sqrt{\kappa}})\cap\mathcal{B}_{\textnormal{bln}}({1}/{2c_{l}})\cap\mathcal{B}_{\mu}$ and $(U_{s},V_{s})\in\mathcal{C}(U_{s-1},V_{s-1},\frac{\sigma_{r}^{*}}{4^{s}c_{e}\kappa})$ for every $1\leq s\leq t+1$ then
@@
We start by proving that $(U_{t+1},V_{t+1})\in\mathcal{B}_{\textnormal{bln}}(\frac{1}{c_{l}})$. By eq.˜40 of Lemma C.8,
@@
By eq.˜51 of Lemma D.1, for large enough $c_{e}$, for each iteration step $0\leq k\leq t$
@@
Inserting eq.˜53 into eq.˜52 gives,
@@
By the assumption $(U_{0},V_{0})\in\mathcal{B}_{\textnormal{err}}\left(\frac{1}{c_{e}\sqrt{\kappa}}\right)$ it follows that $\|U_{0}V_{0}^{\top}-L^{*}\|\leq\frac{\sigma_{r}^{*}}{c_{e}\sqrt{\kappa}}$. Inserting this inequality and eq.˜55 into eq.˜54 gives
@@
To upper bound the first term on the RHS of eq.˜56 we use the following lemma.
@@
To upper bound the RHS of eq.˜57 we employ the following two lemmas. The first lemma upper bounds the first term and the second lemma lower bounds the second term.
@@
To upper bound $\mathcal{L}^{t}_{I_{t}^{\text{in}}}(U_{t+1},V_{t+1})$ we insert eq.˜59 and eq.˜58 into eq.˜57 which yields,
@@
Inserting eq.˜61 into eq.˜60 and using the fact that $\|S\|_{F({\Omega_{t}})}=\|S^{*}\|_{F(I_{t}^{\text{out}})}$ gives,
@@
Inserting eq.˜63 and eq.˜64 into eq.˜62 gives,